OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −(x − 2) y − 2 = −(x + 3)

OpenStudy (anonymous):

use y-y1=m(x-x1) or y-y2=m(x-x2) where m=y2-y1/x2-x1

OpenStudy (anonymous):

i tried i got b

OpenStudy (anonymous):

what is slope? -1/5 or 5